A defibrillator is used during a heart attack to restore the heart to its normal beating pattern. A defibrillator passes 16 A of current through the torso of a person in 2.5 ms. (a) How much charge moves during this time? (b) How many electrons pass through the wires connected to the patient?
Answer #1

Here,

a) for the total charge

total charge = current * time

total charge = 16 * 0.0025

total charge = 0.04 C

total charge moved is 0.04 C

part b)

number of electrons pass = charge/e

number of electrons pass = 0.04/(1.602 *10^-19)

number of electrons pass = 2.5 *10^17

the number of electrons passing is 2.5 *1017